OpenStudy (anonymous):

An arithmetic sequence is represented by the explicit formula A(n) = 2 + 9(n - 1). What is the recursive formula? A. A(n) = A(n - 1) + 2 B. A(n) = A(n - 1) + 9 C. A(n) = A(n - 1) - 9 D. A(n) = A(n - 1) - 2

OpenStudy (zarkon):

\[A(n) = 2 + 9(n - 1)\] \[=2 + 9(n - 2+1)\] \[=2 + 9(n - 2)+9\] \[=A(n-1)+9\]

OpenStudy (anonymous):

tnks